a theatre owner wants to provide enough seating for 973 people. The main floor has 25 rows of 26 seats in each row. If the balcony has 19 rows, how many seats must be in each row to satisfy the owner's seating requirements?
Answer by jim_thompson5910(35256)

The main floor has 25*26 = 650 seats.

So you need 973 - 650 = 323 more seats

There are 19 rows in the balcony and let's say there are x seats in each row. Since you need 323 seats, this means

19x = 323

x = 323/19

x = 17

So there must be 17 seats in each row up in the balcony.